# Optimal Transmit Signal Design for Multi-Target MIMO Sensing Exploiting Prior Information

### v1

@@ VIII Numerical Results
In this section, we present numerical results to evaluate the performance of our proposed transmit signal designs in multi-target sensing. Consider a MIMO radar system with $N_{t}=10$ transmit antennas and $N_{r}=12$ receive antennas under a ULA layout with half-wavelength antenna spacing. The array steering vectors are thus given by $\boldsymbol{a}(\theta_{m})\!=\![e^{\frac{-\!j\pi(N_{t}\!-\!1)\!\sin\!\theta_{m}}{2}},\!e^{\frac{-\!j\pi(N_{t}\!-\!3)\!\sin\!\theta_{m}}{2}},...,\!e^{\frac{j\pi(N_{t}\!-\!1)\!\sin\!\theta_{m}}{2}}]^{T}$ and $\boldsymbol{b}(\theta_{m})\!\!=[e^{\frac{-\!j\pi(N_{r}\!-\!1)\sin\theta_{m}}{2}},\!e^{\frac{-\!j\pi(N_{r}-3)\sin\theta_{m}}{2}},...,e^{\frac{j\pi(N_{r}-1)\sin\theta_{m}}{2}}]^{T}$. Each target’s angle to be sensed lies in the region $[-\frac{\pi}{2},\frac{\pi}{2})$. The number of samples of the transmit probing signal is set as $L\!\!=\!\!25$, and the average noise power at the BS receiver is $\sigma^{2}\!=\!-90$ dBm. Unless otherwise specified, we set $M\!=\!2$ and $P\!=\!30$ dBm. In practice, each target’s angle is typically concentrated around one or multiple nominal angles. Motivated by this, we consider a practical PDF for $\bm{\theta}$, where the angles of different targets are independent of each other, and each angle $\theta_{m}$ has a PDF following the Gaussian mixture model given by $p_{\Theta_{m}}(\theta_{m})=\sum_{i=1}^{K_{m}}{\frac{p_{m,i}}{\sqrt{2\pi}\sigma_{m,i}}e^{-\frac{(\theta_{m}-\bar{\theta}_{m,i})^{2}}{2\sigma_{m,i}^{2}}}},\ \forall m$. $K_{m}$ denotes the number of Gaussian PDF components of the $m$-th target; $p_{m,i}$, $\bar{\theta}_{m,i}\in[-\frac{\pi}{2},\frac{\pi}{2})$, and $\sigma_{m,i}^{2}$ represent the weight, mean, and variance of the $i$-th Gaussian PDF, respectively. We set $K_{1}\!\!=\!\!4$, $\bar{\theta}_{1,1}\!\!=\!\!-0.77$, $\bar{\theta}_{1,2}\!=\!-0.4$, $\bar{\theta}_{1,3}\!=\!-0.14$, $\bar{\theta}_{1,4}\!=\!0.82$, $p_{1,1}\!=\!0.17$, $p_{1,2}\!=\!0.33$, $p_{1,3}\!=\!0.15$, $p_{1,4}\!=\!0.35$, $\sigma_{1,1}^{2}\!\!=\!\!\sigma_{1,2}^{2}\!\!=\!\!\sigma_{1,3}^{2}\!\!=\!\!10^{-2}$, $\sigma_{1,1}^{2}\!\!=\!\!10^{-2.5}$; and $K_{2}\!\!=\!\!4$, $\bar{\theta}_{2,1}\!\!=\!\!-0.73$, $\bar{\theta}_{2,2}\!\!=\!\!-0.45$, $\bar{\theta}_{2,3}\!\!=\!\!0.76$, $\bar{\theta}_{2,4}\!\!=\!\!0.8$, $p_{2,1}\!\!=\!\!0.16$, $p_{2,2}\!\!=\!\!0.39$, $p_{2,3}\!\!=\!\!0.25$, $p_{2,4}\!\!=\!\!0.2$, $\sigma_{2,1}^{2}\!\!=\!\!\sigma_{2,3}^{2}\!\!=\!\!10^{-2.5}$, $\sigma_{2,2}^{2}\!\!=\!\!\sigma_{2,4}^{2}\!\!=\!\!10^{-2}$. We further set $\sigma_{\alpha_{m}}^{2}\!\!=\!\!10^{-8}$ for all $m$, $\xi\!\!=\!\!0.5$, $\varepsilon\!\!=\!\!0.5$, and $\Delta_{n}\!\!=\!\!10^{-3},\forall n$ for the proposed sum-of-ratios iterative algorithm.
@@ IX Conclusions
This paper studied the transmit sample covariance matrix optimization in a MIMO radar system for sensing the unknown and random angles of multiple targets by exploiting their joint distribution. To analytically quantify the multi-target sensing performance, the PCRB for the MSE matrix was derived, which is general for any prior distribution. Then, the transmit covariance matrix optimization problem was formulated to minimize the PCRB for the sum MSE in estimating all targets’ angles. By leveraging Lagrange duality theory, it was unveiled that the optimal transmit covariance matrix has a rank-one structure. Moreover, a novel sum-of-ratios iterative algorithm was proposed which finds the optimal solution with low complexity. Numerical results showed that the proposed algorithm has a fast convergence rate, and outperforms various benchmark schemes in terms of PCRB.

### v2

@@ VIII Numerical Results
In this section, we present numerical results to evaluate the performance of our proposed transmit signal designs in multi-target sensing. Consider a MIMO radar system with $N_{t}=10$ transmit antennas and $N_{r}=12$ receive antennas under a ULA layout with half-wavelength antenna spacing. The array steering vectors are thus given by $\boldsymbol{a}(\theta_{m})\!=\![e^{\frac{-\!j\pi(N_{t}\!-\!1)\!\sin\!\theta_{m}}{2}},\!e^{\frac{-\!j\pi(N_{t}\!-\!3)\!\sin\!\theta_{m}}{2}},...,\!e^{\frac{j\pi(N_{t}\!-\!1)\!\sin\!\theta_{m}}{2}}]^{T}$ and $\boldsymbol{b}(\theta_{m})\!\!=[e^{\frac{-\!j\pi(N_{r}\!-\!1)\sin\theta_{m}}{2}},\!e^{\frac{-\!j\pi(N_{r}-3)\sin\theta_{m}}{2}},...,e^{\frac{j\pi(N_{r}-1)\sin\theta_{m}}{2}}]^{T}$. Each target’s angle to be sensed lies in the region $[-\frac{\pi}{2},\frac{\pi}{2})$. The number of samples of the transmit probing signal is set as $L\!\!=\!\!25$, and the average noise power at the BS receiver is $\sigma^{2}\!=\!-90$ dBm. Unless otherwise specified, we set $M=2$  [^2] It is worth noting that this work is applicable to any value of $M$. and $P\!=\!30$ dBm. In practice, each target’s angle is typically concentrated around one or multiple nominal angles. Motivated by this, we consider a practical PDF for $\bm{\theta}$, where the angles of different targets are independent of each other, and each angle $\theta_{m}$ has a PDF following the Gaussian mixture model given by $p_{\Theta_{m}}(\theta_{m})=\sum_{i=1}^{K_{m}}{\frac{p_{m,i}}{\sqrt{2\pi}\sigma_{m,i}}e^{-\frac{(\theta_{m}-\bar{\theta}_{m,i})^{2}}{2\sigma_{m,i}^{2}}}},\ \forall m$. $K_{m}$ denotes the number of Gaussian PDF components of the $m$-th target; $p_{m,i}$, $\bar{\theta}_{m,i}\in[-\frac{\pi}{2},\frac{\pi}{2})$, and $\sigma_{m,i}^{2}$ represent the weight, mean, and variance of the $i$-th Gaussian PDF, respectively. We set $K_{1}\!\!=\!\!4$, $\bar{\theta}_{1,1}\!\!=\!\!-0.77$, $\bar{\theta}_{1,2}\!=\!-0.4$, $\bar{\theta}_{1,3}\!=\!-0.14$, $\bar{\theta}_{1,4}\!=\!0.82$, $p_{1,1}\!=\!0.17$, $p_{1,2}\!=\!0.33$, $p_{1,3}\!=\!0.15$, $p_{1,4}\!=\!0.35$, $\sigma_{1,1}^{2}\!\!=\!\!\sigma_{1,2}^{2}\!\!=\!\!\sigma_{1,3}^{2}\!\!=\!\!10^{-2}$, $\sigma_{1,1}^{2}\!\!=\!\!10^{-2.5}$; and $K_{2}\!\!=\!\!4$, $\bar{\theta}_{2,1}\!\!=\!\!-0.73$, $\bar{\theta}_{2,2}\!\!=\!\!-0.45$, $\bar{\theta}_{2,3}\!\!=\!\!0.76$, $\bar{\theta}_{2,4}\!\!=\!\!0.8$, $p_{2,1}\!\!=\!\!0.16$, $p_{2,2}\!\!=\!\!0.39$, $p_{2,3}\!\!=\!\!0.25$, $p_{2,4}\!\!=\!\!0.2$, $\sigma_{2,1}^{2}\!\!=\!\!\sigma_{2,3}^{2}\!\!=\!\!10^{-2.5}$, $\sigma_{2,2}^{2}\!\!=\!\!\sigma_{2,4}^{2}\!\!=\!\!10^{-2}$. We further set $\sigma_{\alpha_{m}}^{2}\!\!=\!\!10^{-11}$ for all $m$, $\xi\!\!=\!\!0.5$, $\varepsilon\!\!=\!\!0.5$, and $\Delta_{n}\!\!=\!\!10^{-3},\forall n$ for the proposed sum-of-ratios iterative algorithm.
@@ IX Conclusions
This paper studied the transmit sample covariance matrix optimization in a MIMO radar system for sensing the unknown and random angles of multiple targets by exploiting their joint distribution. To analytically quantify the multi-target sensing performance, the PCRB for the MSE matrix was derived, which is general for any prior distribution. Then, the transmit covariance matrix optimization problem was formulated to minimize the PCRB for the sum MSE in estimating all targets’ angles. Moreover, a novel sum-of-ratios iterative algorithm was proposed which finds the optimal solution with low complexity. Numerical results showed that the proposed algorithm has a fast convergence rate, and outperforms various benchmark schemes in terms of PCRB.